The ratio of two numbers  is 2:3 and their sum is 85.find the number 

Grade:12th pass

1 Answers

Arun
5 years ago
Dear student
 
Let no are x and y
 
x /y = 2 /3
3 x = 2y
 
Now
 
x + y = 85
2 x + 2y = 170
5x = 170
 
x = 34
Hence y = 51
 
Hope it helps
 
 
Regards
Arun